One number is eight times larger than another. Four times the larger number added to two times the smaller one equals 20. What are the two number.
Answer by Alan3354(69443)

One number is eight times larger than another. Four times the larger number added to two times the smaller one equals 20. What are the two number.
------------
B = bigger number
S = smaller
---
B = 9S
--
4B + 2S = 20
4*9S + 2S = 20
38S = 20
S = 10/19
B = 90/19